# WeightedPose: Generalizable Cross-Pose Estimation via Weighted SVD

### v1

@@ Abstract
We present a novel method for robotic manipulation tasks in human environments that require reasoning about the 3D geometric relationship between a pair of objects. Traditional end-to-end trained policies, which map from pixel observations to low-level robot actions, struggle to reason about complex pose relationships and have difficulty generalizing to unseen object configurations. To address these challenges, we propose a method that learns to reason about the 3D geometric relationship between objects, focusing on the relationship between key parts on one object with respect to key parts on another object. Our standalone model utilizes Weighted SVD to reason about both pose relationships between articulated parts and between free-floating objects. This approach allows the robot to understand the relationship between the oven door and the oven body, as well as the relationship between the lasagna plate and the oven, for example. By considering the 3D geometric relationship between objects, our method enables robots to perform complex manipulation tasks that reason about object-centric representations. We open source the code and demonstrate the results here  [^1] https://github.com/harryzhangOG/weighted-pose/tree/main.
@@ 1 Introduction
Many manipulation tasks require a robot to move an object to a location relative to another object. For example, a cooking robot may need to place a lasagna in an oven, place a pot on a stove, place a plate in a microwave, place a mug onto a mug rack, or place a cup onto a shelf. Understanding and placing objects in task-specific locations is a key skill for robots operating in human environments. Further, this skill should generalize to novel objects within the training categories, such as placing new trays into the oven or new mugs onto a mug rack.
A common approach in robot learning is to train a policy “end-to-end," mapping from pixel observations to low-level robot actions. However, end-to-end trained policies cannot easily reason about complex pose relationships such as the ones described above, and they have difficulty generalizing to unseen object configurations. In contrast, we propose a method that learns to reason about the 3D geometric relationship between a pair of objects. For the type of tasks defined above, the robot needs to reason about the relationship between key parts on one object with respect to key parts on another object. Specifically, the robot needs to reason about both the relationship between the oven door and the oven body, as well as the relationship between the lasagna plate and the oven. We propose a standalone model that utilizes Weighted SVD to reason about both pose relationship between articulated parts and between free-floating objects.
@@ 2 Background
Object Pose Estimation: Pose estimation is the task of detecting and inferring the 6DoF pose of an object, which includes its position and orientation, with respect to some previously defined object reference frame (19; 28; 36; 10; 11; 32). Recent work (20; 27; 33; 21) proposed to use 3D semantic keypoints as an alternative form of object representation. While keypoint-based methods can generalize within an object class, they require a significant amount of hand annotated data or access to a simulated version of the task to learn to estimate the keypoint locations. In contrast, our method is able to learn from just 10 real-world demonstrations. Another approach is to use dense embeddings, such as Dense Object Nets (DON) (9) and Neural Descriptor Fields (NDF) (31; 7), which achieve generalization across classes by predicting dense embeddings in the observation and matching them to embeddings of the demonstration objects. However, DON (9) and NDF (31) assume that the target object is moved relative to a static reference object in a “known canonical configuration" (e.g. the pose of the mug rack in NDF (31) is assumed to be known and fixed). In contrast, our method reasons about the geometric relationship between a pair of objects and hence does not need to assume a static environment.
@@
Articulated Object Manipulation: Manipulation of articulated objects and other objects with non-rigid properties remains an open research area due to the objects’ complex geometries and kinematics. Previous work proposed manipulating such objects by hand-designed analytical methods, such as the immobilization of a chain of hinged objects by (5; 29; 1; 2; 39). Berenson et al. (3) proposed a planning framework for manipulation under kinematic constraints. Katz et al. (15) proposed a method to learn such manipulation policies in the real-world using a grounded relational representation learned through interaction. With the development of larger-scale datasets of articulated objects such as the PartNet dataset by Mo et al. (22) and Partnet-Mobility by Xiang et al. (35); Sim et al. (30); Jin et al. (14), several works have proposed learning methods based on large-scale simulation and supervised visual learning. Mo et al. (23) proposed to learn articulation manipulation policies through large-scale simulation and visual affordance learning. Xu et al. (37) proposed a system that learns articulation affordances as well as an action scoring module, which can be used to articulate objects. Several works have focused specifically on visual recognition and estimation of articulation parameters, learning to predict the pose (42; 45; 43; 38; 34; 12; 16) and identify articulation parameters (13; 41; 44) to obtain action trajectories. Moreover, (24; 4; 6; 17; 18; 8) tackle the problem using statistical motion planning.
@@ 3 Method
Point Displacement Loss: Instead of directly supervising the rotation and translation (as is done in DCP), we supervise the predicted transformation using its effect on the points. For this loss, we take the point clouds of the objects in the demonstration configuration, and transform each cloud by a random transform, $\mathbf{\hat{P}}_{\mathcal{A}}=\mathbf{T}_{\alpha}\mathbf{P}_{\mathcal{A}}$, and $\mathbf{\hat{P}}_{\mathcal{B}}=\mathbf{T}_{\beta}\mathbf{P}_{\mathcal{B}}$. This would give us a ground truth transform of $\mathbf{T}_{{\mathcal{A}}{\mathcal{B}}}^{GT}=\mathbf{T}_{\beta}\mathbf{T}_{\alpha}^{-1}$; the inverse of this transform would move object $\mathcal{B}$ to the correct position relative to object $\mathcal{A}$. Using this ground truth transform, we compute the MSE loss between the correctly transformed points and the points transformed using our prediction.
@@
Direct Correspondence Loss. While the Point Displacement Loss best describes errors seen at inference time, it can lead to correspondences that are inaccurate but whose errors average to the correct pose. To improve these errors we directly supervise the learned correspondences $\tilde{V}_{\mathcal{A}}$ and $\tilde{V}_{\mathcal{B}}$:
@@
Correspondence Consistency Loss. Furthermore, a consistency loss can be used. This loss penalizes correspondences that deviate from the final predicted transform. A benefit of this loss is that it can help the network learn to respect the rigidity of the object, while it is still learning to accurately place the object. Note, that this is similar to the Direct Correspondence Loss, but uses the predicted transform as opposed to the ground truth one. As such, this loss requires no ground truth:
@@ 4 Experiments
We define a PartNet-Mobility Placement task as placing a given action object relative to an anchor object based on a semantic goal position. This would require the model to be able to output cross-pose for both the articulated parts and the free-floating objects. We select a set of household furniture objects from the PartNet-Mobility dataset (35) and a set of small rigid objects released with the Ravens simulation environment (40). When the model is required to estimate goal pose for the articulated part, following the terminology in (25), the action object would be the articulated part (e.g. doors, drawers), and the anchor object would be the static body of the furniture. In contrast, when the model is estimating goal pose for free floating objects, then it becomes the same setting as in (25).

### v2

@@ Abstract
We introduce a new approach for robotic manipulation tasks in human settings that necessitates understanding the 3D geometric connections between a pair of objects. Conventional end-to-end training approaches, which convert pixel observations directly into robot actions, often fail to effectively understand complex pose relationships and do not easily adapt to new object configurations. To overcome these issues, our method focuses on learning the 3D geometric relationships, particularly how critical parts of one object relate to those of another. We employ Weighted SVD in our standalone model to analyze pose relationships both in articulated parts and in free-floating objects. For instance, our model can comprehend the spatial relationship between an oven door and the oven body, as well as between a lasagna plate and the oven. By concentrating on the 3D geometric connections, our strategy empowers robots to carry out intricate manipulation tasks based on object-centric perspectives  [^1] https://github.com/harryzhangOG/weighted-pose/tree/main.
@@ 1 Introduction
Numerous robotic tasks involve relocating an object to a position relative to another object. For instance, a culinary robot might be required to put a lasagna in an oven, set a pot on a stove, place a dish in a microwave, position a mug on a rack, or situate a cup on a shelf. Mastery in positioning objects according to task-specific needs is crucial for robots working in human environments. Moreover, this capability should extend to new items within known categories, such as inserting different trays into an oven or different mugs onto a rack.
Traditionally in robot learning, policies are trained end2end to translate pixel data directly into robot actions. However, these end-to-end policies struggle with understanding intricate pose relationships like those mentioned and often fail to adapt to new object configurations. In response, we suggest a method that focuses on understanding the 3D geometric relationships between pairs of objects. For the aforementioned tasks, this involves discerning the spatial relationships between critical parts of one object in relation to another—for example, analyzing both the connection between an oven door and the main oven body and between a lasagna plate and the oven itself. Our proposed independent model uses Weighted SVD to analyze these pose relationships, whether they involve articulated parts or free-floating objects.
@@ 2 Background
Object Pose Estimation: Pose estimation involves detecting and determining the 6DoF (six degrees of freedom) pose of an object, which encompasses its position and orientation relative to a predefined object reference frame. Recent research has suggested using 3D semantic keypoints as a novel object representation form (19; 28; 36; 10; 11; 32). Recent work (20; 27; 33; 21) Although keypoint-based methods show generalization within object classes, they rely heavily on extensive hand-annotated data or simulated environments to learn keypoint locations. In contrast, our method can learn from merely 10 real-world demonstrations. Another technique involves using dense embeddings, such as Dense Object Nets (DON) and Neural Descriptor Fields (NDF), which facilitate class-wide generalization by predicting and matching dense embeddings from observations to those of demonstration objects. In contrast, our method is able to learn from just 10 real-world demonstrations. Another approach is to use dense embeddings, such as Dense Object Nets (DON) (9) and Neural Descriptor Fields (NDF) (31; 7), which achieve generalization across classes by predicting dense embeddings in the observation and matching them to embeddings of the demonstration objects. However, both DON and NDF presuppose that the target object is manipulated relative to a static reference in a "known canonical configuration" (for example, the fixed pose of a mug rack in NDF). Unlike these methods, our approach understands the geometric relationships between pairs of objects, allowing it to function without assuming a static environment (31). In contrast, our method reasons about the geometric relationship between a pair of objects and hence does not need to assume a static environment.
@@
Articulated Object Manipulation: The manipulation of articulated objects and those with non-rigid characteristics continues to be a challenging area of research, largely because of their intricate geometries and kinematics. Earlier research has suggested handling these objects using hand-crafted analytical methods, for example, by immobilizing a sequence of hinged objects through (5; 29; 1; 2; 39). Berenson et al. (3) proposed a planning framework for manipulation under kinematic constraints. Katz et al. (15) proposed an approach to acquire manipulation policies in real-world settings by using a relational representation that is grounded and learned through interaction. This method was enhanced by the creation of expansive datasets of articulated objects like the PartNet dataset Mo et al. (22) and Partnet-Mobility by Xiang et al. (35); Sim et al. (30); Jin et al. (14), everal studies have suggested using learning techniques that combine large-scale simulation with supervised visual learning. Mo et al. (23)introduced an approach to develop articulation manipulation policies through extensive simulation and visual affordance learning. Xu et al. (37) suggested a framework that acquires knowledge of articulation affordances and includes an action scoring module, which is utilized to manipulate objects. Various studies have specifically concentrated on the visual identification and calculation of articulation parameters, developing methods to estimate the pose (42; 45; 43; 38; 34; 12; 16) and identify articulation parameters (13; 41; 44) to obtain action trajectories. Moreover, (24; 4; 6; 17; 18; 8) tackle the problem using statistical motion planning.
@@ 3 Method
Point Displacement Loss:
Rather than directly supervising the rotation and translation (as seen in DCP), we manage the predicted transformation by observing its impact on the points. To do this, we utilize the point clouds from the objects in their demonstration setup and apply a random transformation to each cloud for this loss, $\mathbf{\hat{P}}_{\mathcal{A}}=\mathbf{T}_{\alpha}\mathbf{P}_{\mathcal{A}}$, and $\mathbf{\hat{P}}_{\mathcal{B}}=\mathbf{T}_{\beta}\mathbf{P}_{\mathcal{B}}$. This would give us a ground truth transform of $\mathbf{T}_{{\mathcal{A}}{\mathcal{B}}}^{GT}=\mathbf{T}_{\beta}\mathbf{T}_{\alpha}^{-1}$; the inverse of this transform would move object $\mathcal{B}$ to the correct position relative to object $\mathcal{A}$. Using the actual transformation as a reference, we calculate the Mean Squared Error (MSE) loss by comparing the correctly transformed points to those transformed based on our predicted values.
@@
Direct Correspondence Loss. Although the Point Displacement Loss accurately reflects errors observed during inference, it may result in correspondences that are individually inaccurate yet average out to the correct pose. To enhance these errors, we directly supervise the correspondences that are learned $\tilde{V}_{\mathcal{A}}$ and $\tilde{V}_{\mathcal{B}}$:
@@
Correspondence Consistency Loss. Additionally, a consistency loss may be employed. This loss penalizes any discrepancies between correspondences and the final predicted transformation. An advantage of this loss is its ability to encourage the network to maintain object rigidity while it learns precise object placement. It is important to note that this is akin to the Direct Correspondence Loss, except that it relies on the predicted transformation rather than the ground truth. Therefore, this loss does not require any ground truth data:
@@ 4 Experiments
We describe a PartNet-Mobility Placement task as the act of positioning an action object in relation to an anchor object according to a predefined semantic goal position. This task demands that the model generate a cross-pose for both articulated parts and free-floating objects. We have chosen a collection of household furniture items from the PartNet-Mobility dataset for this purpose (35) and a set of small rigid objects released with the Ravens simulation environment (40). When the model is required to estimate goal pose for the articulated part, following the terminology in (25), articulated part serves as the action object, while the static body of the furniture acts as the anchor object. Conversely, when the model estimates the goal pose for free-floating objects, the scenario aligns with the same setting as previously discussed (25).